square root of x = square root of 4x-12
square both sides =
x = 4x-12
[Or you could write it like this, I don't know what you guys can understand...]
^(x) = ^(4x-12)
^(x)*2 = ^(4x-12)*2
x = 4x-12

This is all I can figure out by way of my text book. Do you know the next steps in the solution. If so, thank you very much! I'm sorry for bothering you, and I hope you can understand the problem clearly enough!
~Nuit D'Automne~

Answer by queenofit(120)

If I understand your work so far you have an equation that had the square root on both sides of an equal sign. You squared both sides to get rid of the square root sign, which is the correct thing to do. you were left with the equation x=4x-12. You have little confidence in your abilities because you have done this correctly and have come to the easy part. Just solve for x now.
x=4x-12
x-4x=-12
-3x=-12
x=4
I suppose this answer could be plus or minus 4 since you had a square root to begin with and when you have a sqare root you really do not know if that number is a positive or a negative. However, I believe you did this problem correctly and the answer is 4. Good job...